OpenStudy (anonymous):

Write the equation y - 2 = 2(x - 3) in standard form

jimthompson5910 (jim_thompson5910):

y - 2 = 2(x - 3) y - 2 = 2x-6 y = 2x-6+2 y = 2x-4 ---------------------------- y-2x = -4 -2x+y = -4 2x-y = 4